OpenStudy (anonymous):

What is the slope of a line that passes through the point (-2, 3) and is parallel to a line that passes through (3, 7) and (-2, -8)?

OpenStudy (anonymous):

parallel slopes are equal.\[m = (-8-7)/(-2-3) = -15/-5 = 3\]

OpenStudy (anonymous):

y - 7 = 3 (x - 3) y = 3x -9 + 7 y = 3x -2

OpenStudy (anonymous):

so is the answer 3 ? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I also wrote the equaation of hte parallel line :)